# Fingerprinting Implementations of Cryptographic Primitives and Protocols that Use Post-Quantum Algorithms

### v3

@@ Abstract
Fingerprinting is a technique used to create profiles of systems to identify potential threats, weaknesses, and malicious activities. For cryptographic primitives and network protocols such fingerprinting can be exploited by attackers to conduct denial-of-service, key recovery, or downgrade attacks.
In this paper, we study the feasibility of fingerprinting post-quantum (PQ) algorithms by analyzing implementations of primitives such as key exchange and digital signatures, their usage within secure protocols, and integration into SNARK generation libraries. Unlike traditional cryptographic algorithms, PQ algorithms have larger memory needs and higher computation costs. Our research examines liboqs and CIRCL libraries, PQ implementations of the TLS, SSH, QUIC, OpenVPN and Open ID Connect(OIDC) protocols across Windows, Ubuntu, and MacOS, and two SNARK generation and verification libraries - pysnark  and lattice_zksnark  on Ubuntu. Experimental results reveal that, depending on computational workload, (1) we can distinguish between between classical and PQ key exchange and digital signature algorithms with accuracies of 98% and 100%, respectively; (2) we can guess the used algorithm within PQ key exchange or digital signature algorithms with 97% and 86% accuracy, respectively; (3) we can distinguish between implementation of the same algorithm in the liboqs and CIRCL libraries for key exchange and digital signature with 97% and 100% accuracy, respectively; and (4) within the same library –CIRCL  we can distinguish the PQ and hybrid key exchange algorithm implementations with 97% accuracy. In the case of secure protocols, it is possible to (1) discern whether the key exchange is classical or PO, and (2) identify the specific PQ key exchange algorithm employed. Finally, the SNARK generation and verification of pysnark and lattice_zksnark  are distinguishable with an accuracy of 100%. We demonstrate the applicability of our fingerprinting methods to real systems by applying them to the Tranco dataset to identify domains that use PQ key exchange for TLS, and by integrating them in QUARTZ, an open source risk and threat analyzer created by CISCO.
@@ 1 Introduction
Quantum computing is a field of information technology that uses the principles of quantum mechanics to process and manipulate data in fundamentally new ways, allowing it to solve complex problems exponentially faster than classical computers. One of the fields where quantum computers will have a significant impact is cryptography as cryptographic primitives such as RSA digital signatures and Diffie-Hellman key exchange rely their security on computational problems. While quantum computers are not yet practical for most real-world attacks on classical cryptography, a new research direction emerged to design schemes that can withstand such attacks and secure data against future quantum attackers.
NIST conducted a multi-year process to solicit, evaluate, and standardize cryptographic algorithms that are secure against attacks by quantum computers for encryption and digital signatures [53] leading to the development of post-quantum cryptography (PQC). The results of the competition were announced in 2023, drafts for comments were publicly available for encryption [58] and digital signatures [59, 60], and the first PQC standards  [57]: FIPS 203 [54], FIPS 204 [55], and FIPS 205 [56] were announced in August 2024. NIST is also working on standards and recommendations for "how to transition to PQC" [52].
@@
Implementations of PQC primitives have been publicly made available in libraries and integrated in secure communication protocols and secure storage. Examples of libraries include – liboqs [62], CIRCL [26], AWS-LC [6], and pqm4[70]. The liboqs  library is a project of Open Quantum Safe, part of the Linux Foundation’s Post-Quantum Cryptography Alliance and provides a broad platform for integrating and evaluating post-quantum algorithms, in general-purpose operating systems like Linux, macOS, and Windows. The CIRCL  library developed by Cloudflare, offers post-quantum and other advanced cryptographic algorithms with a focus on the programming language Go. AWS-LC (AWS-LibCrypto) is Amazon’s cryptographic library, which has started integrating post-quantum algorithms to enhance security within AWS services. Finally, the pqm4  library is an academic research project and is optimized for the ARM Cortex-M4 microcontroller, focusing on post-quantum algorithms suited for resource-constrained embedded systems.
Additionally, it has become increasingly important to safeguard secure protocols such as QUIC, SSH, DTLS [44], TLS, OAuth2.0 against the potential threats posed by quantum computing, Google, Microsoft, Cloudflare, Amazon have started to incorporate quantum algorithms along with classical ones in their implementations of aforementioned protocols.
Post-quantum algorithms are characterized by distinct memory footprints, often demanding greater memory resources due to the intricate nature of their operations and data structures. Furthermore, post quantum algorithms have different CPU cycle counts, which stems from the use of mathematical operations involving significantly larger keys and ciphertext sizes. These attributes render post-quantum algorithms noticeably distinct from classical cryptographic methods, creating opportunity for fingerprinting.
Fingerprinting is a technique used to create profiles of systems to identify potential threats, weaknesses, and the presence of malicious activities. For cryptographic primitives and network protocols such fingerprinting can be exploited by attackers to conduct denial-of-service, key recovery, or downgrade attacks. It can also be used for risk assessment to evaluate the vulnerability to attacks. Previous research has explored methods for fingerprinting secure protocols using classical cryptography. The study in [79] demonstrated how the network flow of OpenVPN could be fingerprinted by identifying fixed patterns within OpenVPN traffic. Similarly, the authors in [66] presented an attack on Tor, revealing that traffic analysis could be used to identify visited websites, despite the use of encryption, by exploiting unique traffic patterns generated by different websites. Additional studies [78, 77] have shown that machine learning models applied to encrypted traffic can effectively identify the underlying protocols, highlighting vulnerabilities in encrypted communications.
@@
Our contribution. In this work we explore if it is possible to fingerprint post quantum primitives from their library implementations, their usage within protocols that protect Internet communication, or more complex cryptographic constructions such as SNARK-s widely used in blockchains. We start with publicly available post quantum libraries from Microsoft, Cloudflare, Amazon to show that, when compared to classical algorithms in terms of computational cost and memory overhead their post quantum counterparts are easily distinguishable by using machine learning models. We collected data from liboqs and CIRCL running on Windows, Ubuntu, and MacOS under a scenario when other intensive computation tasks share the resources with the post quantum libraries. We formulate distinguishing between key exchange and digital signature as classifying problems and use machine learning algorithms to solve the problem. Ensemble learning models, particularly XGBoost, exhibit the highest performance across various classification tasks. Specifically, it achieved 98% accuracy in classifying classical vs. post quantum key exchange algorithms and 100% accuracy for classical vs. post-quantum digital signatures. For post-quantum key exchange algorithms, the classification accuracy reached 97%, while for post-quantum digital signatures, it was 86%. XGBoost also effectively distinguished between implementations of the same algorithm in the liboqs and CIRCL libraries, achieving 96% accuracy in key exchange and 100% in digital signature. For post-quantum and hybrid key exchange classification within CIRCL library, it achieved 97% accuracy.
@@
We next proceed to see if fingerprinting is possible when post quantum algorithms are implemented for key exchanges in widely used protocols such as TLS[35], SSH[32], QUIC[39], OpenVPN [65] and Open ID Connect(OIDC) [33, 34]. We start by gathering packet captures for these protocols with Wireshark and tcpdump. For classical packet captures we either use publicly available captures from Cloudflare[73], Wireshark [1] or use tcpdump on existing classical implementations of those protocols to generate packet captures locally in our devices. For protocols using TLS, such as OIDC, QUIC, and OpenVPN, we isolate key exchange packets and then compare the key sizes within the TLS layer to determine whether the connection employs classical or post quantum algorithm. Similarly, for SSH, we isolate the packets and identify key exchange from information embedded in the SSH layer instead of TLS. By integrating these approaches, we successfully identify the key exchange algorithms in the majority of TLS 1.3 and all SSH connections with the exception of a small number of connections such ascertain Windows TLS connections using custom headers and fragmented TLS 1.2 connections from OpenVPN.
@@
Finally, we study the feasibility of fingerprinting SNARK generation and verification in the pysnark  (classical) and lattice_zksnark  (post-quantum) libraries. XGBoost is the most effective, achieving 100% accuracy between the post-quantum and classical library implementations.
@@
We demonstrate the applicability of our fingerprinting methods to real systems by applying them to the Tranco dataset to identify domains that use PQ key exchange for TLS, and by integrating them in QUARTZ, an open source risk and threat analyzer created by CISCO.
@@ 2.1 KEM and Digital Signature Algorithms
The algorithms are categorized based on the mathematical problems they rely on, each offering different strengths, weaknesses, and applications. A summary of the algorithms is provided in Table 1 and we expound on the categories below.
Lattice-based cryptography is one of the most promising and versatile areas of post-quantum cryptography. It relies on problems in lattice theory, such as the Shortest Vector Problem (SVP) and the Learning With Errors (LWE) problem. These problems are believed to be difficult for both classical and quantum computers to solve. Lattice-based cryptographic schemes are valued for their efficiency in computation and storage, making them suitable for a broad range of applications, including key exchange, encryption, and digital signatures. Key exchange algorithms CRYSTALS-Kyber[14], NTRU[31], SABER[21] and digital signature scheme CRYSTALS-Dilithium are finalists in the NIST post-quantum cryptography standardisation process.
Code-based cryptography is another well-established category, rooted in the hardness of decoding random linear codes—a problem considered difficult even for quantum computers. While this category of algorithms offer small ciphertexts and fast encryption/decryption, their main drawback is the large size of public keys, which can be challenging to manage. Key exchange algorithm Classic McEliece[45] is a finalist in the NIST post-quantum cryptography standardization process for this category.
@@
Multivariate polynomial cryptography is based on the difficulty of solving systems of multivariate quadratic equations over finite fields, a problem known to be NP-hard. Algorithms in this category, such as Rainbow[22] and GeMSS[15], are primarily used for digital signatures. These schemes provide a valuable alternative for post-quantum security, despite their complexity and inefficiency due to large keys and signatures. Rainbow is a finalist in the NIST post-quantum cryptography standardization process for this category.
Isogeny-based cryptography represents a more specialized area, using the mathematical structure of elliptic curves and the difficulty of computing isogenies between them. Schemes like SIKE (Supersingular Isogeny Key Encapsulation) [40] are compact and have relatively small key sizes, which is a significant advantage for certain applications like secure communications in constrained environments. However, in July 2022, KU Leuven researchers broke the SIKE algorithm, a NIST fourth-round candidate, using a classical computer in 62 minutes, highlighting vulnerabilities in its underlying supersingular isogeny problem.
@@
Standardization. Earlier in 2022, as part of the standardization efforts, NIST selected four algorithms: CRYSTALS-Kyber, CRYSTALS-Dilithium, Sphincs + and FALCON for standardization, with draft versions of three standards released in 2023. In August 2024, NIST released the first PQC standards  [57]: FIPS 203 [54], FIPS 204 [55], and FIPS 205 [56]. NIST is working to standardize a second set of algorithms. The fourth draft standard, based on FALCON, was expected in late 2024. But as of 2025, there has been no announcement of a final release yet.
@@
Lattice-based CRYSTALS-Kyber Kyber512, Kyber768, Kyber1024 Key Encapsulation liboqs[62], CIRCL[26], pqm4[70], AWS-LC [6]
FRODOKem[13] FRODO-640, FRODO-976, FRODO-1344 Key Encapsulation liboqs, CIRCL
@@
FALCON [28] FALCON-512, FALCON-1024 Digital Signature liboqs, pqm4
ML-DSA[51] ML-DSA-44, ML-DSA-65, ML-DSA-87 Digital Signature liboqs
@@
BIKE [8] BIKE-L1, BIKE-L3, BIKE-L5 Key Encapsulation liboqs, pqm4
HQC[5] HQC-128, HQC-192, HQC-256 Key Encapsulation liboqs
Hash-based SPHINCS+ [9] SPHINCS+-SHA2-128-simple, SPHINCS+-SHA2-192-simple, SPHINCS+- SHA2-256-simple Digital Signature liboqs
XMSS[17] XMSS, XMSS-MT Digital Signature N/A
@@
Zero Knowledge-based PICNIC[18] PICNIC2-L1-FS, PICNIC2-L3-FS, PICNIC2-L5-FS Digital Signature N/A
Table 1: Post-Quantum Cryptography Algorithms

### v4

# Classifying Implementations of Cryptographic Primitives and Protocols that Use Post-Quantum Algorithms

@@ Abstract
Classification techniques can be used to analyze system behaviors, network protocols, and cryptographic primitives based on identifiable traits. While useful for defensive purposes, such classification can also be leveraged by attackers to infer system configurations, detect vulnerabilities, and tailor specific attacks — including denial-of-service, key recovery, or downgrade attacks — against the categorized targets.
In this paper, we study the feasibility of classifying post-quantum (PQ) algorithms by analyzing implementations of primitives such as key exchange and digital signatures, their usage within secure protocols, and integration into SNARK generation libraries. Unlike traditional cryptographic algorithms, PQ algorithms have larger memory needs and variable computation costs. Our research examines two post-quantum cryptography libraries, liboqs  and CIRCL. We evaluate their implementations of TLS, SSH, QUIC, OpenVPN, and OpenID Connect (OIDC) across Windows, Ubuntu, and macOS. Additionally, we analyze two SNARK generation and verification libraries—pysnark  and lattice_zksnark—on Ubuntu. Experimental results reveal that, depending on computational workload, (1) we can distinguish between classical and PQ key exchange and digital signature algorithms with accuracies of 98% and 100%, respectively; (2) we can determine the used algorithm within PQ key exchange or digital signature algorithms with 97% and 86% accuracy, respectively; (3) we can distinguish between implementations of the same algorithm in the liboqs and CIRCL libraries for key exchange and digital signature with 97% and 100% accuracy, respectively; and (4) within the same library –CIRCL  we can distinguish the PQ and hybrid key exchange algorithm implementations with 97% accuracy. In the case of secure protocols, it is possible to (1) discern whether the key exchange is classical or PQ, and (2) identify the specific PQ key exchange algorithm employed. Finally, the SNARK generation and verification of pysnark and lattice_zksnark  are distinguishable with an accuracy of 100%. We demonstrate the applicability of our classification methods to real systems by applying them to the Tranco dataset to identify domains that use PQ key exchange for TLS, and by integrating them in QUARTZ, an open source risk and threat analyzer created by CISCO.
@@ 1 Introduction
Quantum computing leverages quantum mechanics to process data in fundamentally new ways, enabling exponentially faster solutions to complex problems than classical computers. A major area of impact is cryptography, as schemes like RSA and Diffie-Hellman rely on computational problems vulnerable to quantum attacks. Although practical quantum computers are not yet available, researchers are developing cryptographic methods resilient to future quantum threats and designed to prevent “harvest-now, decrypt-later” attacks.
NIST conducted a multi-year initiative to evaluate and standardize cryptographic algorithms resistant to quantum attacks, leading to the establishment of post-quantum cryptography (PQC) [nist-pqc-details]. The process concluded in 2023 with public drafts for encryption [nist-fips203_encryption] and digital signatures [nist-fips204_ds, nist-fips205_hds], followed by the release of the first PQC standards—FIPS 203 [NIST-pqc-standard-fips203], FIPS 204 [NIST-pqc-standard-fips204], and FIPS 205 [NIST-pqc-standard-fips205]—in August 2024 [NIST-pqc-standard-Aug132024]. In March 2025, NIST selected Hamming Quasi-Cyclic (HQC) as a fifth key encapsulation mechanism, providing a non-lattice alternative to ML-KEM (FIPS 203), with a draft standard expected in 2026 and finalization by 2027. Concurrently, draft NIST IR 8547, Transition to Post-Quantum Cryptography Standards, offers guidance on migration planning, risk assessment, crypto agility, and asset inventory, recommending deprecation of quantum-vulnerable algorithms (RSA, ECDSA, EdDSA, DH, ECDH) by 2030 and their full retirement by 2035.
@@
Implementations of post-quantum cryptographic (PQC) primitives are now publicly available in libraries and integrated into secure communication protocols and storage systems. Notable examples include liboqs [liboqs], CIRCL [circl], AWS-LC [aws-lc], pqm4 [pqm4], PQClean [pqclean], and libpqcrypto [libpqcrypto]. liboqs, developed by the Open Quantum Safe project under the Linux Foundation’s Post-Quantum Cryptography Alliance, supports integration and evaluation of PQC algorithms across major operating systems. CIRCL, by Cloudflare, provides advanced cryptographic algorithms for the Go programming language. Amazon’s AWS-LC library incorporates PQC schemes to strengthen AWS security. The pqm4 library targets resource-constrained embedded systems on ARM Cortex-M4. PQClean offers clean, portable C implementations of NIST PQC candidates for reference and reuse, while libpqcrypto, from the PQCRYPTO project, collects diverse encryption and signature schemes for research and benchmarking.
Additionally, it has become crucial to secure protocols like QUIC, SSH, DTLS [rfc9147], TLS, and OAuth 2.0 against quantum threats. Major tech companies are already adopting hybrid and quantum-safe solutions: Google, Microsoft, Cloudflare, and Amazon integrate quantum algorithms into their protocols; IBM offers quantum-safe TLS; Meta uses hybrid key exchange; Akamai prepares PQC deployment across its CDN; and Cisco embeds quantum-safe methods in its networking products—signaling an industry-wide shift toward quantum-resilient infrastructure.
Post-quantum algorithms have distinct memory footprints and CPU cycle patterns, due to their complex operations, larger keys, and ciphertexts. These characteristics make them identifiable through classification techniques that analyze behavioral patterns and protocol-specific responses. Such classification can expose systems to risks, as attackers may exploit identifiable traits to launch denial-of-service, key recovery, or downgrade attacks. Conversely, it can aid defenders in assessing risks and identifying vulnerable components. Previous research has focused on classifying secure protocols using classical cryptography, underscoring the growing importance of similar approaches for post-quantum systems.
Our contribution. In this work we investigate whether post-quantum cryptographic primitives can be classified based on their library implementations, use in secure communication protocols, or integration in advanced constructions like SNARKs. The classification task is challenging because post-quantum algorithms exhibit diverse computational and memory characteristics—some requiring more resources, others less, and some closely matching classical algorithms. These inconsistencies make rule-based approaches unreliable, motivating the use of machine learning models to capture subtle behavioral differences between classical and post-quantum schemes.
@@
We start with publicly available post quantum libraries from Microsoft, Cloudflare, Amazon to show that, when compared to classical algorithms in terms of computational cost and memory overhead their post quantum counterparts are easily distinguishable by using machine learning models. We collected data from liboqs and CIRCL running on Windows, Ubuntu, and MacOS under a scenario when other intensive computation tasks share the resources with the post quantum libraries. We formulate distinguishing key exchange schemes as well as signature schemes, employing machine learning methods to carry out the classification. Ensemble learning models, particularly XGBoost, exhibit the highest performance across various classification tasks. Specifically, it achieved 98% accuracy in classifying classical vs. post quantum key exchange algorithms and 100% accuracy for classical vs. post-quantum digital signatures. For post-quantum key exchange algorithms, the classification accuracy reached 97%, while for post-quantum digital signatures, it was 86%. XGBoost also effectively distinguished between implementations of the same algorithm in the liboqs and CIRCL libraries, achieving 96% accuracy in key exchange and 100% in digital signature. For post-quantum and hybrid key exchange classification within CIRCL library, it achieved 97% accuracy.
@@
We next proceed to see if classification is possible when post quantum algorithms are implemented for key exchanges in widely used protocols such as TLS[tls], SSH[rfc4253], QUIC[rfc9000], OpenVPN [openvpn] and Open ID Connect(OIDC) [rfc6749, rfc6750]. We start by gathering packet captures for these protocols with Wireshark and tcpdump. For classical packet captures we either use publicly available captures from Cloudflare[qacafe_sample_captures], Wireshark [wireshark_sample_captures] or use tcpdump on existing classical implementations of those protocols to generate packet captures locally in our devices. For protocols using TLS, such as OIDC, QUIC, and OpenVPN, we isolate the key exchange packets and then compare the key sizes within the TLS layer to determine whether the connection employs a classical or post quantum algorithm. Similarly, for SSH, we isolate the packets and identify key exchange from information embedded in the SSH layer instead of TLS. By integrating these approaches, we successfully identify the key exchange algorithms in the majority of TLS 1.3 and all SSH connections with the exception of a small number of connections such as certain Windows TLS connections using custom headers and fragmented TLS 1.2 connections from OpenVPN.
@@
Finally, we study the feasibility of classifying SNARK generation and verification in the pysnark  (classical) and lattice_zksnark  (post-quantum) libraries. XGBoost is the most effective, achieving 100% accuracy between the post-quantum and classical library implementations.
@@
We demonstrate the applicability of our classification methods to real systems by applying them to the Tranco dataset to identify domains that use PQ key exchange for TLS, and by integrating them in QUARTZ, an open source risk and threat analyzer created by CISCO.
@@
Availability. Our code is available at https://anonymous.4open.science/r/Fingerprint-809D/README.md.
@@ 2.1 KEM and Digital Signature Algorithms
The algorithms are categorized based on the mathematical problems they rely on, each offering different strengths, weaknesses, and applications. A summary of the algorithms is provided in Table I and we expound on the categories below.
Lattice-based cryptography is one of the most promising and versatile areas of post-quantum cryptography. It relies on problems in lattice theory, such as the Shortest Vector Problem (SVP) and the Learning With Errors (LWE) problem. These problems are believed to be difficult for both classical and quantum computers to solve. Lattice-based cryptographic schemes are valued for their efficiency in computation and storage, making them suitable for a broad range of applications, including key exchange, encryption, and digital signatures. Key exchange algorithms CRYSTALS-Kyber[kyber2018], NTRU[ntru1998], SABER[saber2018] and digital signature scheme CRYSTALS-Dilithium are finalists in the NIST post-quantum cryptography standardization process.
Code-based cryptography is rooted in the hardness of decoding random linear codes—a problem considered difficult even for quantum computers. While this category of algorithms offer small ciphertexts and fast encryption/decryption, their main drawback is the large size of public keys, which can be challenging to manage. Key exchange algorithm Classic McEliece[mceliece1978] is a finalist in the NIST post-quantum cryptography standardization process for this category.
@@
Multivariate polynomial cryptography is based on the difficulty of solving systems of multivariate quadratic equations over finite fields, a problem known to be NP-hard. Algorithms in this category, such as Rainbow[rainbow2005] and GeMSS[gemss2018], are primarily used for digital signatures. These schemes provide a valuable alternative for post-quantum security, despite their complexity and inefficiency due to large keys and signatures. Rainbow is a finalist in the NIST post-quantum cryptography standardization process for this category.
Isogeny-based cryptography represents a more specialized area, using the mathematical structure of elliptic curves and the difficulty of computing isogenies between them. Schemes like SIKE (Supersingular Isogeny Key Encapsulation) [sike2017] are compact and have relatively small key sizes, which is a significant advantage for certain applications like secure communications in constrained environments. However, in July 2022, KU Leuven researchers broke the SIKE algorithm, a NIST fourth-round candidate, using a classical computer in 62 minutes, highlighting vulnerabilities in its underlying supersingular isogeny problem.
@@
Standardization. Earlier in 2022, NIST selected four algorithms: CRYSTALS-Kyber, CRYSTALS-Dilithium, Sphincs + and FALCON for standardization, with draft versions of three standards released in 2023. In August 2024, NIST released the first PQC standards  [NIST-pqc-standard-Aug132024]: FIPS 203 [NIST-pqc-standard-fips203], FIPS 204 [NIST-pqc-standard-fips204], and FIPS 205 [NIST-pqc-standard-fips205]. NIST is working to standardize a second set of algorithms. The fourth draft standard, based on FALCON, was expected in late 2024. But as of 2025, there has been no announcement of a final release yet.
@@
Lattice-based CRYSTALS-Kyber Kyber512, Kyber768, Kyber1024 Key Encapsulation liboqs[liboqs], CIRCL[circl], pqm4[pqm4], AWS-LC [aws-lc]
FRODOKem[frodo2016] FRODO-640, FRODO-976, FRODO-1344 Key Encapsulation liboqs, CIRCL
@@
FALCON [FALCON2018] FALCON-512, FALCON-1024 Digital Signature liboqs, pqm4
ML-DSA[ml-dsa] ML-DSA-44, ML-DSA-65, ML-DSA-87 Digital Signature liboqs
@@
BIKE [bike2021] BIKE-L1, BIKE-L3, BIKE-L5 Key Encapsulation liboqs, pqm4
HQC[hqc2018] HQC-128, HQC-192, HQC-256 Key Encapsulation liboqs
Hash-based SPHINCS+ [sphincs2019] SPHINCS+-SHA2-128-simple, SPHINCS+-SHA2-192-simple, SPHINCS+- SHA2-256-simple Digital Signature liboqs
XMSS[xmss2011] XMSS, XMSS-MT Digital Signature N/A
@@
Zero Knowledge-based PICNIC[picnic2019] PICNIC2-L1-FS, PICNIC2-L3-FS, PICNIC2-L5-FS Digital Signature N/A
TABLE I: Post-Quantum Cryptography Algorithms